OpenStudy (anonymous):

A team needs to save up at least $400 for new uniforms and currently has $210. They decide to hold a fundraiser where they can earn $7 in profits for each item sold. What inequality could be solved to determine the minimum number, n, of items to be sold? would it be greater than or greater than or equal to?

OpenStudy (anonymous):

can any one solve this shy1980 A team needs to save up at least $400 for new uniforms and currently has $210. They decide to hold a fundraiser where they can earn $7 in profits for each item sold. What inequality could be solved to determine the minimum number, n, of items to be sold? would it be greater than or greater than or equal to?

OpenStudy (anonymous):

A team needs to save up $400 for new uniforms and currently has $210. They decide to hold a fundraiser where they can earn $7 in profits for each item sold. What equation could be solved to determine the number, n, of items to be sold? . 400 = 7n + 210 (this is what they're looking for) solving for n: 190 = 7n 190/7 items = n I think there is an error on This question cause :p Cause I think they currently has 190 and the answer would be 210/3=30 items :D

OpenStudy (anonymous):

okay I see what you are saying but it wants us to write it in a number form and I thought it would be 210 + 7n < 400 just dont know if this is right any thoughs

OpenStudy (texaschic101):

210 + 7n >= 400

OpenStudy (texaschic101):

it has to be at least 400....so equal or greater then 400